OpenStudy (anonymous):

Find seven, positive, all different, whole numbers x1,x2,x3,x4,x5,x6,x7 such that 1/x1+1/x2+1/x3+1/x4+1/x5+1/x6+1/x7=1

OpenStudy (anonymous):

i got the answer by intution, but mathematics doesnt accept it so plz gimme a procedure for such kindaa problems

OpenStudy (anonymous):

i got the question online there solution is also there i want a different one coz it looks ugly. and i thought the question is good so I put it up

OpenStudy (anonymous):

no there is an exact answer i just want the procedure man. if u want i'll tell u how i got to the answer. but its no procedure

OpenStudy (anonymous):

How about \[ 1= \frac 12 + \frac 12 \] and then \[ 1= \frac 13 + \frac 1 6 + \frac 12 \] and then something like this ...

OpenStudy (anonymous):

ya i got the same way but i want a procedure say if i am asked to split x among y rational nubers what should i do?

OpenStudy (anonymous):

Btw the interesting question is "how many ways we can write 1 as a sum of X unique reciprocal numbers?"

OpenStudy (anonymous):

what's wrong using this there too?

OpenStudy (anonymous):

i dunno i get this feeling that , this is no procedure. just an arbit guess which is fruitful